# Diffusion Posterior Sampling is Computationally Intractable

### v1

@@ 1 Introduction
Over the past few years, diffusion models have emerged as a powerful way for representing distributions of images. Such models, such as Dall-E [${}^{+}$22] and Stable Diffusion [${}^{+}$21], are very effective at learning and sampling from distributions. These models can then be used as priors for a wide variety of downstream tasks, including inpainting, superresolution, and MRI reconstruction.
Diffusion models are based on representing the smoothed scores of the desired distribution. For a distribution $p(x)$, we define the smoothed distribution $p_{\sigma}(x)$ to be $p$ convolved with $\mathcal{N}(0,\sigma^{2}I)$. These have corresponding smoothed scores $s_{\sigma}(x):=\nabla\log p_{\sigma}(x)$. Given the smoothed scores, the distribution $p$ can be sampled using an SDE [HJA20] or an ODE [SME21]. Moreover, the smoothed score is the minimizer of what is known as the score-matching objective, which can be estimated from samples.
Sampling via diffusion models is fairly well understood from a theoretical perspective. The sampling SDE and ODE are both fast (polynomial time) and robust (tolerating $L_{2}$ error in the estimation of the smoothed score). Moreover, with polynomial training samples of the distribution, the empirical risk minimizer (ERM) of the score matching objective will have bounded $L_{2}$ error, leading to accurate samples [BMR20, GPPX23]. So diffusion models give fast and robust unconditional samples.
@@
Posterior sampling has many appealing properties for image reconstruction tasks. For example, if you want to identify $x$ precisely, posterior sampling is within a factor 2 of the minimum error possible for every measurement model and every error metric [${}^{+}$21]. When ambiguities do arise, posterior sampling has appealing fairness guarantees with respect to sensitive attributes [${}^{+}$21].
Given the appeal of posterior sampling, the natural question is: is efficient posterior sampling possible given approximate smoothed scores? A large number of recent papers [${}^{+}$21, ${}^{+}$23, KVE21, ${}^{+}$23, SVMK23, KEES22, DS24] have studied algorithms for posterior sampling, with promising empirical results. But all these fail on some inputs; can we find a better posterior sampling algorithm that is fast and robust in all cases?
@@
$\operatorname{\mathbb{P}}[\left\|x-\widehat{x}\right\|>\tau]<\delta.$
@@
The theory of diffusion models [${}^{+}$23] says that the diffusion process gives an unconditional sampler for well-modeled distributions that takes polynomial time (with the precise polynomial improved by subsequent work [BBDD24]).
@@
Suppose that one-way functions exist. Then for any $d^{0.01}<m<d/2$, there exists a $10$-well-modeled distribution over $\mathbb{R}^{d}$, and linear measurement model with $m$ measurements and noise parameter $\beta=\Theta(\frac{1}{\log^{2}d})$, such that $(\frac{1}{10},\frac{1}{10})$-posterior sampling requires superpolynomial time.
To be a one-way function, inversion must take superpolynomial time (on average). For many one-way function candidates used in cryptography, the best known inversion algorithms actually take exponential time. Under the stronger assumption that some one-way function exists that requires exponential time to invert with non-negligible probability, we can show that posterior sampling takes $2^{\Omega(m)}$ time:
@@
For any $m\leq d/2$, suppose that there exists a one-way function $f:\{\pm 1\}^{m}\to\{\pm 1\}^{m}$ that requires $2^{\Omega(m)}$ time to invert. Then for any $C>1$, there exists a $C$-well-modeled distribution over $\mathbb{R}^{d}$ and linear measurement model with $m$ measurements and noise level $\beta=\frac{1}{C^{2}\log^{2}d}$, such that $(\frac{1}{10},\frac{1}{10})$-posterior sampling takes at least $2^{\Omega(m)}$ time.
@@
The lower bound produces a “well-modeled” distribution, meaning that the scores are representable by a polynomial-size neural network, but there is no requirement that the network be shallow. One could instead consider only shallow networks; the same theorem holds, except that $f$ must also be computable by a shallow depth network. Many candidate one-way functions can be computed in $\textup{{NC}}^{0}$ (i.e., by a constant-depth circuit) [AIK04], so the cryptographic assumption is still mild.
@@ 2 Related Work
Diffusion models [SDWMG15, DN21, SE19] have emerged as the most popular approach to deep generative modeling of images, serving as the backbone for the recent impressive results in text-to-image generation [${}^{+}$22, ${}^{+}$21], along with state-of-the-art results in video [${}^{+}$23, ${}^{+}$22] and audio [${}^{+}$21, ${}^{+}$21] generation.
Noisy linear inverse problems capture a broad class of applications such as image inpainting, super-resolution, MRI reconstruction, deblurring, and denoising. The empirical success of diffusion models has motivated their use as a data prior for linear inverse problems, without task-specific training. There have been several recent theoretical and empirical works [${}^{+}$21, ${}^{+}$23, KVE21, ${}^{+}$23, SVMK23, KEES22, DS24] proposing algorithms to sample from the posterior of a noisy linear measurement. We highlight some of these approaches below.
@@
One class of approaches [${}^{+}$23, KVE21, SVMK23] approximates the intractable posterior score $\nabla\log p_{t}(x_{t}|y)=\nabla\log p_{t}(x_{t})+\nabla\log p(y|x_{t})$ at time $t$ of the reverse diffusion process, and uses this approximation to sample. Here, $y=Ax_{0}+\eta$ is the noisy measurement of $x_{0}\sim p_{0}$, where $p_{t}$ is the density at time $t$. For instance, [${}^{+}$23] proposes the approximation $\nabla\log p(y|x_{t})\approx\nabla\log p(x|\operatorname*{\mathbb{E}}\left[x_{0}|x_{t}\right])$, thereby incurring error quantified by the so-called Jensen gap. [SVMK23] proposes an approximation based on the pseudoinverse of $A$, while [KVE21] proposes to use the score of the posterior wrt measurement $y_{t}$ of $x_{t}$.
@@
Another approach, first introduced in the context of inpainting [${}^{+}$22], replaces the observed coordinates of the sample with a noisy version of the observation during the reverse diffusion process. An extension was proposed for general noisy linear measurements [KEES22]. This approach essentially also attempts to sample from an approximation to the posterior.
@@
A recent set of works [${}^{+}$23, ${}^{+}$23, DS24] makes use of Sequential Monte Carlo (SMC) methods to sample from the posterior. These methods are guaranteed to sample from the correct distribution as the number of particles goes to $\infty$. Our paper implies a lower bound on the number of particles necessary for good convergence. Assuming one-way functions exist, polynomially many particles are insufficient in general, so that these algorithms takes superpolynomial time; assuming some one-way function requires exponential time to invert, particle filtering requires exponentially many particles for convergence.
@@ 3 Proof Overview – Lower Bound
Given a one-way function $f:\{-1,1\}^{d}\to\{-1,1\}^{d}$, consider the distribution that is uniform over $(s,f(s))\in\{-1,1\}^{2d}$ for all $s\in\{-1,1\}^{d}$. This distribution is easy to sample from unconditionally: sample $s$ uniformly, then compute $f(s)$. At the same time, posterior sampling is hard: if you observe the last $d$ bits, i.e. $f(s)$, a posterior sample should be from $f^{-1}(f(s))$; and if $f$ is a one-way function, finding any point in this support is computationally intractable on average.
@@ 3.1 Lower Bound Instance
For any $b\in\{-1,1\}$, let $\psi_{b}$ be the density of a standard Gaussian discretized to multiples of $\varepsilon$, with phase either 00 or $\frac{\varepsilon}{2}$ depending on $b$:
@@
$x_{i}$ $\sim w_{1}(x_{i}-R\cdot s_{i})$ for $\text{for }i\leq d$
$x_{i}$ $\sim\psi_{f(s)_{i-d}}$ for $\text{for }i>d.$
@@
The measurement process then takes sample $x\sim\widetilde{g}$ and computes $Ax+\eta$, where $\eta=N(0,\beta^{2}I_{d^{\prime}})$ and $A=\begin{pmatrix}0^{d^{\prime}\times d}&I_{d^{\prime}}\end{pmatrix}.$ That is, we observe the last $d^{\prime}$ bits of $x$, with variance $\beta^{2}$ Gaussian noise added to each coordinate.
@@ 3.2 Posterior Sampling Implies Inversion
$\operatorname{\mathbb{P}}_{s,\mathsf{A}}[f(\mathsf{A}(f(s)))\neq f(s)]\leq\frac{3}{4}$
Take some $z\in\{\pm 1\}^{d^{\prime}}$. Our goal is to compute $f^{-1}(z)$, using the posterior sampler for $\widetilde{g}$. To do this, we take a sample $\overline{z}_{i}\sim\psi_{z_{i}}*N(0,\beta^{2})$, for $i\in[d^{\prime}]$, and feed in $\overline{z}$ into our posterior sampler, to output $\widehat{x}$. We then take the first $d$ bits of $\widehat{x}$, round each entry to the nearest ±$\pm 1$, and output the result.
To see why this works, let’s analyze what the resulting conditional distribution looks like. First, note that any sample $x\sim\widetilde{g}$ encodes some $(s,f(s))$ coordinate-wise so that the encoding of $f(s)$ is one of two discretizations of a normal distribution, with width $\varepsilon$, offset by $\varepsilon/2$ from each other (see Figure 1). Furthermore, since $\beta\ll\varepsilon$, these two encodings are distinguishable with high probability even after adding noise with variance $\beta^{2}$. Therefore, with high probability, our sample $\overline{z}$, which is a noised and discretized encoding of the input $z$ we want to invert, will be such that each coordinate is within $\varepsilon/4$ of the correct discretization. Consequently, a posterior sample with this observation will correspond to an encoding of $(s,f(s))$ where $s=f^{-1}(z)$, with high probability. The first $d$ bits of this encoding are just the bits of $f^{-1}(z)$ smoothed by a gaussian with variance $1/R^{2}$, and since $R\gg 1$, rounding these coordinates to the nearest ±$\pm 1$ returns $f^{-1}(z)$, with high probability.
@@
Suppose $d^{0.01}\leq m\leq d/2$ and one-way functions exist. Then, for $\widetilde{g}$ as defined in Definition 3.2 with $\varepsilon=\frac{1}{C\sqrt{\log d}}$ and $R=C\log d$, and linear measurement model with noise parameter $\beta=\frac{1}{C^{2}\log^{2}d}$ and measurement matrix $A\in\mathbb{R}^{m\times d}$, $(\frac{1}{10},\frac{1}{10})$-posterior sampling takes superpolynomial time.
@@ 3.3.1 ReLU Approximation for Score of Product Distribution
$\operatorname*{\mathbb{E}}\left[(s_{\sigma}(x)-s_{\sigma}(\alpha_{x}))^{2}\right]$ $\lesssim\gamma^{2}\sigma^{4}\operatorname*{\mathbb{E}}[\sup_{c}s_{\sigma}^{\prime}(x+c)^{2}]$
$\lesssim\gamma^{2}$
@@ 3.3.2 ReLU Approximation for Large $\sigma$
Figure 3: $\psi_{1}$ and $\psi_{-1}$ are discretized Gaussians with discretization width $\varepsilon$ and phase 00 and $\varepsilon/2$ respectively. If we convolve with $\mathcal{N}(0,\sigma^{2})$, we get a distribution close to Gaussian when $\sigma\geq\varepsilon$ for each of $\psi_{1},\psi_{-1}$.
@@ 3.3.3 ReLU Approximation for Small $\sigma$
where $h_{s}$ is the product distribution that is Gaussian with mean $R\cdot s_{i}$ in the first $d$ coordinates and is a smoothed discretized Gaussian with mean 00 in the remaining $d^{\prime}$ coordinates.
@@ 4 Proof Overview - Upper Bound
$\operatorname*{\mathbb{E}}_{x}\left[e^{-\frac{\left\|Ax-y\right\|^{2}}{2\beta^{2}}}\right]\geq\operatorname{\mathbb{P}}_{x}\left[\left\|Ax-y\right\|\leq O(\beta\sqrt{m})\right]\cdot e^{-O(m)}.$
@@
$\operatorname{\mathbb{P}}_{x}\left[\left\|Ax-y\right\|\leq O(\beta)\right]\geq\beta.$
For well-modeled distributions, the covariance matrix of $x$ has constant singular values. Then with arbitrarily high probability, $x$ is $O(1)$ in each direction. Since every singular value of $A$ is at most 1, the projection $Ax$ onto $\mathbb{R}$ will lie in $[-C,+C]$ for some constant $C$ with arbitrarily high probability.
We divide $[-C,+C]$ into $N=\frac{2C}{\beta}$ segments of length $\beta$, forming set $S$. Now we only need to prove that with arbitrarily high probability over $y$, there exists a segment $\theta\in S$ satisfying for all $x\in\theta$, $|x-y|\leq O(\beta)$ , and $\operatorname{\mathbb{P}}_{x\sim{\mathcal{D}_{x}}}[Ax\in\theta]\gtrsim\beta.$ For any constant $c>0$, define
$S^{\prime}:=\{\theta\in S\mid\operatorname{\mathbb{P}}_{x\sim{\mathcal{D}_{x}}}[Ax\in\theta]>\frac{c}{N}\}.$
@@
$\operatorname{\mathbb{P}}_{x}\left[\left\|Ax-y\right\|\leq O(\beta\sqrt{m})\right]\geq\Theta(\beta)^{m}.$
@@
For $(x,\widehat{x})$ drawn from this coupling, we know from our previous analysis that rejection sampling based on $x$ is correct. But the algorithm only knows $\widehat{x}$, which changes its behavior in two ways: (1) it chooses to accept based on $p(y\mid\widehat{x})$ rather than $p(y\mid x)$, and (2) it returns $\widehat{x}$ rather than $x$ on acceptance. The perturbation from (2) is easily within our tolerance, since it is $\frac{1}{d^{2C}}$ close to $x$ with arbitrarily high probability.

### v2

@@ 1 Introduction
Over the past few years, diffusion models have emerged as a powerful way for representing distributions of images. Such models, such as Dall-E ramesh2022hierarchical and Stable Diffusion rombach2022highresolution, are very effective at learning and sampling from distributions. These models can then be used as priors for a wide variety of downstream tasks, including inpainting, superresolution, and MRI reconstruction.
Diffusion models are based on representing the smoothed scores of the desired distribution. For a distribution $p(x)$, we define the smoothed distribution $p_{\sigma}(x)$ to be $p$ convolved with $\mathcal{N}(0,\sigma^{2}I)$. These have corresponding smoothed scores $s_{\sigma}(x):=\nabla\log p_{\sigma}(x)$. Given the smoothed scores, the distribution $p$ can be sampled using an SDE ho2020denoising or an ODE song2021denoising. Moreover, the smoothed score is the minimizer of what is known as the score-matching objective, which can be estimated from samples.
Sampling via diffusion models is fairly well understood from a theoretical perspective. The sampling SDE and ODE are both fast (polynomial time) and robust (tolerating $L_{2}$ error in the estimation of the smoothed score). Moreover, with polynomial training samples of the distribution, the empirical risk minimizer (ERM) of the score matching objective will have bounded $L_{2}$ error, leading to accurate samples Block2020GenerativeMW; gupta2023sampleefficient. So diffusion models give fast and robust unconditional samples.
@@
Posterior sampling has many appealing properties for image reconstruction tasks. For example, if you want to identify $x$ precisely, posterior sampling is within a factor 2 of the minimum error possible for every measurement model and every error metric mri_jalal. When ambiguities do arise, posterior sampling has appealing fairness guarantees with respect to sensitive attributes Jalal2021FairnessFI.
Given the appeal of posterior sampling, the natural question is: is efficient posterior sampling possible given approximate smoothed scores? A large number of recent papers mri_jalal; chung2023diffusion; kawar2021snips; trippe2023practical; song2023pseudoinverseguided; kawar2022denoising; particle2024diffusion have studied algorithms for posterior sampling, with promising empirical results. But all these fail on some inputs; can we find a better posterior sampling algorithm that is fast and robust in all cases?
@@
$\operatorname*{\mathbf{Pr}}[\left\|x-\widehat{x}\right\|>\tau]<\delta.$
@@
The theory of diffusion models chen2023sampling says that the diffusion process gives an unconditional sampler for well-modeled distributions that takes polynomial time (with the precise polynomial improved by subsequent work benton2024nearly).
@@
Suppose that one-way functions exist. Then for any $m>d^{0.01}$, there exists a $10$-well-modeled distribution over $\mathbb{R}^{d}$, and linear measurement model with $m$ measurements and noise parameter $\beta=\Theta(\frac{1}{\log^{2}d})$, such that $(\frac{1}{10},\frac{1}{10})$-posterior sampling requires superpolynomial time in $d$.
To be a one-way function, inversion must take superpolynomial time on average. It is widely believed, including for problems based on lattices Lattice and elliptic curves zhandry2019magic, that many one-way function candidates need exponential time to invert. Under the stronger assumption that there exist some one-way functions that require exponential time to invert with non-negligible probability, we can show that posterior sampling takes $2^{\Omega(m)}$ time:
@@
Suppose that there exist one-way functions $f:\{\pm 1\}^{m}\to\{\pm 1\}^{m}$ that require $2^{\Omega(m)}$ time to invert. Then for any $m\leq O(d)$ and $C>1$, there exists a $C$-well-modeled distribution over $\mathbb{R}^{d}$ and linear measurement model with $m$ measurements and noise level $\beta=\frac{1}{C^{2}\log^{2}d}$, such that $(\frac{1}{10},\frac{1}{10})$-posterior sampling takes at least $2^{\Omega(m)}$ time.
@@
The lower bound produces a “well-modeled” distribution, meaning that the scores are representable by a polynomial-size neural network, but there is no requirement that the network be shallow. One could instead consider only shallow networks; the same theorem holds, except that $f$ must also be computable by a shallow depth network. Many candidate one-way functions can be computed in $\textup{{NC}}^{0}$ (i.e., by a constant-depth circuit) NC0, so the cryptographic assumption is still mild.
@@ 2 Related Work
Diffusion models pmlr-v37-sohl-dickstein15; Dhariwal2021DiffusionMB; Song2019GenerativeMB have emerged as the most popular approach to deep generative modeling of images, serving as the backbone for the recent impressive results in text-to-image generation ramesh2022hierarchical; rombach2022highresolution, along with state-of-the-art results in video blattmann2023stable; ho2022video and audio kong2021diffwave; chen2021wavegrad generation.
Noisy linear inverse problems capture a broad class of applications such as image inpainting, super-resolution, MRI reconstruction, deblurring, and denoising. The empirical success of diffusion models has motivated their use as a data prior for linear inverse problems, without task-specific training. There have been several recent theoretical and empirical works mri_jalal; chung2023diffusion; kawar2021snips; trippe2023practical; song2023pseudoinverseguided; kawar2022denoising; particle2024diffusion proposing algorithms to sample from the posterior of a noisy linear measurement. We highlight some of these approaches below.
@@
[…]
@@
Another approach, first introduced in the context of inpainting Lugmayr2022RePaintIU, replaces the observed coordinates of the sample with a noisy version of the observation during the reverse diffusion process. An extension was proposed for general noisy linear measurements kawar2022denoising. This approach essentially also attempts to sample from an approximation to the posterior.
@@
A recent set of works trippe2023practical; trippe2023diffusion; particle2024diffusion makes use of Sequential Monte Carlo (SMC) methods to sample from the posterior. These methods are guaranteed to sample from the correct distribution as the number of particles goes to $\infty$. Our paper implies a lower bound on the number of particles necessary for good convergence. Assuming one-way functions exist, polynomially many particles are insufficient in general, so that these algorithms takes superpolynomial time; assuming some one-way function requires exponential time to invert, particle filtering requires exponentially many particles for convergence.
@@ 3 Proof Overview – Lower Bound
The core idea of our proof is that any general posterior sampler would imply an algorithm that can invert a one-way function. A one-way function is formally defined as follows:
Definition 3.1.
A polynomial-time computable function $f:\{-1,1\}^{*}\to\{-1,1\}^{*}$ is one-way if, for any polynomial-time randomized algorithm $\mathcal{A}$, any constant $c>0$, and all sufficiently large $n$,
$\operatorname*{\mathbf{Pr}}_{x\sim\mathcal{U}_{n}}\left[f(\mathcal{A}(f(x)))=f(x)\right]<n^{-c}$
where $\mathcal{U}_{n}$ is the uniform distribution over $\{-1,1\}^{n}$.
The function $f$ is defined on inputs of arbitrary length; for inputs of length $n$ it can be assumed to have some fixed polynomial output length $m(n)$.
@@
Suppose we have a one-way function $f:\{-1,1\}^{d}\to\{-1,1\}^{d}$, and consider the distribution that is uniform over $(s,f(s))\in\{-1,1\}^{2d}$ for all $s\in\{-1,1\}^{d}$. This distribution is easy to sample from unconditionally: sample $s$ uniformly, then compute $f(s)$. At the same time, posterior sampling is hard: if you observe the last $d$ bits, i.e. $f(s)$, a posterior sample should be from $f^{-1}(f(s))$; and if $f$ is a one-way function, finding any point in this support is computationally intractable on average.
@@ 3.1 Lower Bound Instance
For any $b\in\{-1,1\}$, let $\psi_{b}$ be the density of a standard Gaussian discretized to multiples of $\varepsilon$, with phase either 0 or $\frac{\varepsilon}{2}$ depending on $b$:
@@
$x_{i}$ $\sim w_{1}(x_{i}-R\cdot s_{i})$ $\text{for }i\leq d$
$x_{i}$ $\sim\psi_{f(s)_{i-d}}$ $\text{for }i>d.$
@@
The measurement process then takes sample $x\sim\widetilde{g}$ and computes $Ax+\eta$, where $\eta=\mathcal{N}(0,\beta^{2}I_{d^{\prime}})$ and $A=\begin{pmatrix}0^{d^{\prime}\times d}&I_{d^{\prime}}\end{pmatrix}$. That is, we observe the last $d^{\prime}$ bits of $x$, with variance $\beta^{2}$ Gaussian noise added to each coordinate.
@@ 3.2 Posterior Sampling Implies Inversion
$\operatorname*{\mathbf{Pr}}_{s,\mathsf{A}}[f(\mathsf{A}(f(s)))\neq f(s)]\leq\frac{3}{4}$
Take some $z\in\{\pm 1\}^{d^{\prime}}$. Our goal is to compute $f^{-1}(z)$, using the posterior sampler for $\widetilde{g}$. To do this, we take a sample $\overline{z}_{i}\sim\psi_{z_{i}}*\mathcal{N}(0,\beta^{2})$ for $i\in\{1,\dots,d^{\prime}\}$, and feed in $\overline{z}$ into our posterior sampler, to output $\widehat{x}$. We then take the first $d$ bits of $\widehat{x}$, round each entry to the nearest $\pm 1$, and output the result.
To see why this works, let’s analyze what the resulting conditional distribution looks like. First, note that any sample $x\sim\widetilde{g}$ encodes some $(s,f(s))$ coordinate-wise so that the encoding of $f(s)$ is one of two discretizations of a normal distribution, with width $\varepsilon$, offset by $\varepsilon/2$ from each other (see Figure 1). Furthermore, since $\beta\ll\varepsilon$, these two encodings are distinguishable with high probability even after adding noise with variance $\beta^{2}$. Therefore, with high probability, our sample $\overline{z}$, which is a noised and discretized encoding of the input $z$ we want to invert, will be such that each coordinate is within $\varepsilon/4$ of the correct discretization. Consequently, a posterior sample with this observation will correspond to an encoding of $(s,f(s))$ where $s=f^{-1}(z)$, with high probability. The first $d$ bits of this encoding are just the bits of $f^{-1}(z)$ smoothed by a gaussian with variance $1/R^{2}$, and since $R\gg 1$, rounding these coordinates to the nearest $\pm 1$ returns $f^{-1}(z)$, with high probability.
@@
Suppose $m\geq d^{0.01}$ and one-way functions exist. Then, for $\widetilde{g}$ as defined in Definition 3.3 with $\varepsilon=\frac{1}{C\sqrt{\log d}}$ and $R=C\log d$, and linear measurement model with noise parameter $\beta=\frac{1}{C^{2}\log^{2}d}$ and measurement matrix $A\in\mathbb{R}^{m\times d}$, $(\frac{1}{10},\frac{1}{10})$-posterior sampling takes superpolynomial time.
@@ 3.3.1 ReLU Approximation for Score of Product Distribution
$\operatorname*{\mathbb{E}}\left[(s_{\sigma}(x)-s_{\sigma}(\alpha_{x}))^{2}\right]$ $\lesssim\gamma^{2}\sigma^{4}\operatorname*{\mathbb{E}}[\sup_{c}s_{\sigma}^{\prime}(x+c)^{2}]\lesssim\gamma^{2}$
@@ 3.3.2 ReLU Approximation for Large $\sigma$
Figure 3: $\psi_{1}$ and $\psi_{-1}$ are discretized Gaussians with discretization width $\varepsilon$ and phase 0 and $\varepsilon/2$ respectively. If we convolve with $\mathcal{N}(0,\sigma^{2})$, we get a distribution close to Gaussian when $\sigma\geq\varepsilon$ for each of $\psi_{1},\psi_{-1}$.
@@ 3.3.3 ReLU Approximation for Small $\sigma$
where $h_{s}$ is the product distribution that is Gaussian with mean $R\cdot s_{i}$ in the first $d$ coordinates and is a smoothed discretized Gaussian with mean 0 in the remaining $d^{\prime}$ coordinates.
@@ 4 Proof Overview - Upper Bound
$\operatorname*{\mathbb{E}}_{x}\left[e^{-\frac{\left\|Ax-y\right\|^{2}}{2\beta^{2}}}\right]\geq\operatorname*{\mathbf{Pr}}_{x}\left[\left\|Ax-y\right\|\leq O(\beta\sqrt{m})\right]\cdot e^{-O(m)}.$
@@
$\operatorname*{\mathbf{Pr}}_{x}\left[\left\|Ax-y\right\|\leq O(\beta)\right]\geq\beta.$
For well-modeled distributions, the covariance matrix of $x$ has constant singular values. Then with arbitrarily high probability, $x$ is $O(1)$ in each direction. Since every singular value of $A$ is at most 1, the projection $Ax$ onto $\mathbb{R}$ will lie in [C,+C][-C,+C] for some constant $C$ with arbitrarily high probability.
We divide [C,+C][-C,+C] into $N=\frac{2C}{\beta}$ segments of length $\beta$, forming set $S$. Now we only need to prove that with arbitrarily high probability over $y$, there exists a segment $\theta\in S$ satisfying for all $x\in\theta$, $|x-y|\leq O(\beta)$ , and $\operatorname*{\mathbf{Pr}}_{x\sim{\mathcal{D}_{x}}}[Ax\in\theta]\gtrsim\beta.$ For any constant $c>0$, define
$S^{\prime}:=\{\theta\in S\mid\operatorname*{\mathbf{Pr}}_{x\sim{\mathcal{D}_{x}}}[Ax\in\theta]>\frac{c}{N}\}.$
@@
$\operatorname*{\mathbf{Pr}}_{x}\left[\left\|Ax-y\right\|\leq O(\beta\sqrt{m})\right]\geq\Theta(\beta)^{m}.$
@@
[…]